OpenStudy (liizzyliizz):

y=6cosx^2 (derivative)

OpenStudy (liizzyliizz):

is it -6sinx^2 ?

OpenStudy (amistre64):

paranthesis help, and no

OpenStudy (amistre64):

x^2 pops out a derivative as well

OpenStudy (anonymous):

pop

OpenStudy (anonymous):

not quite - its -12x sinx^2

OpenStudy (amistre64):

it gives birth!! lol

OpenStudy (liizzyliizz):

today is not a good day, I thought i knew this then i take like 3 steps back -_-

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

the chain rule is always in effect

OpenStudy (amistre64):

should be called the pops out rule

OpenStudy (anonymous):

use chain rule lizzy

OpenStudy (liizzyliizz):

wait can u go back and explain how you got the answer my friends and I are a bit lost, we all have different answers.. o_o

OpenStudy (amistre64):

6(f(x)) 6 (sin(g(x))) 6 (sin( (h(x))^2) ) 6 (sin( (x)^2) )

OpenStudy (amistre64):

start inside and work out x' = 1 (x)^2 ' = 2(x) cos'(x^2) = -sin(x^2) .... and of course theres the 6 as a constant to begin with .... 1*2x*-sin(x^2)*6 = -12x sin(x^2)

OpenStudy (amistre64):

gotta read thru the typos

OpenStudy (liizzyliizz):

I see what i was doing wrong i forgot to find the derivative for x^2. but i have another question now for y= 6cos^2x I know u use the chain rule, can i get some help on that one?

OpenStudy (amistre64):